OpenStudy (anonymous):

The area of a circle whose diameter is 18.6 in. is ______ in^2. Round to the hundredths.

OpenStudy (swissgirl):

Area of a circle = pi* radius^2 Now the radius is half the diameter so we would have to divide the diameter by 2 to get the radius So 18.6/2=9.3 Area of circle \(= \pi*9.3^2=271.72\)

OpenStudy (swissgirl):

and there is your answer